OpenStudy (anonymous):

can somebody help me find the general antiderivative P(t=)πtᶺ3+4t

OpenStudy (datanewb):

\[P(t) = \pi t^3 + 4t\] π is just a constant in front of the t^3 term. So is '4' in front of the t term. You can take the antiderivate of t^3 and t separately and then add them together to get your answer. As an example, do you know how to find the antiderivative of \[F(t) =\pi t^2\] That would look like: \[\int\pi t^2 dt = \frac{\pi t^3}{3} + C\] Does that example help you solve your question, @orange123?

OpenStudy (anonymous):

yes it does. Thank you very much datanewb!

OpenStudy (datanewb):

Your welcome.